OpenStudy (anonymous):

lim (h->0) (sqrt(25+h)-5)/h I don't really understand this problem. Can someone help me out?

OpenStudy (anonymous):

draw it, please?

OpenStudy (ash2326):

@hsjunior if you just put h=0, you'll get the form 0/0 so we need to use limits.

OpenStudy (ash2326):

\[\lim_{h\to 0} \frac{\sqrt{25+h}-5}{h}\] Let's rationalize the numerator. Can you try? @hsjumior

OpenStudy (anonymous):

would I have to use l'hopital's theorem?

OpenStudy (ash2326):

Not required, multiply and divide by \[\sqrt{25+h}+5\]

OpenStudy (anonymous):

so is it just multiplying by its conjugate?

OpenStudy (ash2326):

Yes. multiplying and dividing

OpenStudy (anonymous):

does this apply to all problems of this type?

OpenStudy (anonymous):

u know L-Hospitals Rule ?

OpenStudy (ash2326):

On multiplying and dividing you'd get \[\lim_{h \to 0}\frac{(\sqrt{25+h})^2-25}{h(\sqrt{25+h}+5)}\] Now you'll get \[\lim_{h \to 0}\frac{25+h-25}{h(\sqrt{25+h}+5)}\] Can you simplify this?

OpenStudy (anonymous):

\[\frac{ \sqrt{25+h}-5 }{ h }\times \frac{ \sqrt{25+h}+5 }{ \sqrt{25+5}+5 }\] So is it like this?

OpenStudy (ash2326):

yes :)

OpenStudy (anonymous):

and the h then becomes by itself in the numerator, correct?

OpenStudy (anonymous):

Recognize this as the limit definition of the derivative of sqrt(x) at x=25.

OpenStudy (ash2326):

yes @hsjunior

OpenStudy (ash2326):

@Xavier Absolutely

OpenStudy (anonymous):

So if you were to encounter this problem, what would you do first?

OpenStudy (ash2326):

I'd rationalize it first, multiplying and dividing by the conjugate

OpenStudy (anonymous):

so how is it \[\frac{ 1 }{ 10 }\]

OpenStudy (ash2326):

You have \[\lim_{h \to 0}\frac{25+h-25}{h(\sqrt{25+h}+5)}\] Numerator pops out as h \[\lim_{h \to 0}\frac{h}{h(\sqrt{25+h}+5)}\] \[\lim_{h \to 0}\frac{\cancel h}{\cancel h(\sqrt{25+h}+5)}\] We get \[\lim_{h\to 0}\frac{1}{\sqrt{25+h}+5}\] put h=0 \[\frac{1}{\sqrt{25+0}+5}=\frac 1 {5+5}\]

OpenStudy (anonymous):

I understand it now. Thanks guys (:

OpenStudy (ash2326):

Welcome :)

OpenStudy (anonymous):

Do you know all subjects? Is it okay if I ask you a IB Chemistry problem?

OpenStudy (ash2326):

I don't know Chemistry much. You should post the question in Chemistry Subject http://openstudy.com/study#/groups/Chemistry

OpenStudy (anonymous):

okay thanks